OpenStudy (anonymous):

Given that F(x)=sin x, G(x)= cos x, and h(x)=tan x, find the exact value of the composite function please! G(f^-1(4/5))

OpenStudy (jdoe0001):

\(\bf f(x)=sin(x)\qquad g(x)= cos(x)\qquad h(x)=tan(x)\\ \quad \\ \quad \\ f^{-1}(x)\implies sin^{-1}(x)\qquad f^{-1}\left(\frac{4}{5}\right)\implies sin^{-1}\left(\frac{4}{5}\right)\\ \quad \\ g \left(\quad f^{-1}\left(\frac{4}{5}\right)\quad\right)\implies cos\left(sin^{-1}\left(\frac{4}{5}\right)\right)\) same as before, use the pythagorean theorem

OpenStudy (anonymous):

Got it thanks the opposite side ends up being 3 and the cos value is 3/5